OpenStudy (anonymous):

If a triangle has as an area of 144 ft2 and a height of 24 ft, then what is the base?

Parth (parthkohli):

\[\text{Area of a triangle} = {1 \over 2}bh \]\[\text{Here, }144 = {1 \over 2}b\times 24 \]

Parth (parthkohli):

Now let's try to simplify it.\[144 = 0.5 \times b \times 24 \]\[144 = (0.5 \times 24) \times b \]\[ 144 = 12b\]\[ b = {144 \over 12} = \cdots\]

OpenStudy (mayankdevnani):

answer is 12....

Parth (parthkohli):

@peachez Do you get the original idea?

OpenStudy (mayankdevnani):

@peachez

OpenStudy (mayankdevnani):

get idea from parth kohli